# Monocular Facial Appearance Capture in the Wild

### v1

@@ 1 Introduction
Like facial geometry reconstruction, the field of skin appearance estimation is also well-studied in controlled studio environments, where accurate appearance maps can be reconstructed from large setups that obtain multi-view or multi-shot images under calibrated lighting. Following the geometry trend, current research aims to allow facial appearance capture in less constrained settings, for example outdoors using the sun as a single point light [57]. Unfortunately, these methods often make simplifying assumptions, and thus there still exists a large gap in reconstruction quality between current in-the-wild methods and production-ready studio-based capture.
In this work we present a new method for facial appearance capture in the wild, surpassing the level of fidelity of existing lightweight methods. Our approach requires only a short video sequence of a simple head rotation, captured from a single camera in any environment, including indoors or outdoors, on a sunny day or in shadow. Our approach is built on traditional inverse rendering optimization, where a fast differentiable renderer is used to solve for the geometry and appearance parameters together with the environment lighting simultaneously. Different from previous methods, we do not make any assumptions on the lighting condition (e.g. we do not require a sun in the sky), and we explicitly take visibility into account, effectively removing baked-in shading by correctly modeling self-occlusion in our appearance solver. We do so by proposing a novel shading model that combines ray tracing and a visibility modulated pre-filtered environment map. As we will show, our approach leads to more faithful recovery of the appearance properties than existing techniques in the wild. The result is a detailed geometry mesh with textures for diffuse albedo, specular intensity and roughness, which approach the quality achievable from studio-based setups but are captured more flexibly in the wild. As a particular application, our approach allows fast capture of actors on a film set, with resulting assets that can be used directly in traditional VFX pipelines. In summary, we make the following key contributions:
@@
  • A novel shading model that explicitly handles self-occlusion by combining ray tracing with pre-filtered environment maps, achieving high-quality appearance reconstruction from monocular input.
@@ 2 Related Work
Inverse rendering is the process of decomposing the scene into 3D shape, material and illumination by simulating the rendering process and comparing the results against captured images. It has been a popular research topic with the recent advances in novel view synthesis using neural implicit representations [44, 45, 56], 3D Gaussian splatting [35], mesh-based differentiable renderers (rasterizers [36] and path tracers [38, 29]) and using diffusion models as prior [41, 21, 39, 60]. While some existing techniques target a very challenging scenario where the lighting can differ across different images [9, 10, 20], we restrict our discussion here to methods that assume a static unknown environment lighting. PhySG [65] utilizes spherical Gaussians to approximately and efficiently evaluate the rendering equation in closed form. Munkberg et al. [46] propose an efficient end-to-end framework for joint learning of topology, triangle meshes and materials, achieving much faster training and inference compared to previous NeRF-based factorization methods [67, 8, 55]. They also introduce a differentiable formulation of the split-sum approximation of environment lighting to efficiently recover all-frequency lighting. Follow-up work [28] shows that material and lighting decomposition can be further improved with a more realistic shading model, incorporating ray tracing and Monte Carlo integration. Recently, 3D Gaussian Splatting techniques were used in conjunction with physical-based rendering to allow for scene relighting [40, 23, 6, 69, 59].
@@
Traditional face capture studios often employ a multi-view setup with controlled and calibrated lighting conditions to reconstruct the skin appearance [52, 25, 26, 15]. Similar setups were used for facial appearance decomposition with neural and gaussian primitives [53, 64, 62, 61, 1]. Lighter alternative setups were also explored by Lattas et al. [37] and Choi et al. [14] but they still require the subject to be seating in a dedicated space. Recently, the research community has investigated more lightweight setups which are easily accessible to everyone. However, most of these techniques still pose some constraints on the capture environment. CoRA [27] reconstructs relightable 3D face assets from a single co-located smartphone flashlight sequence captured in a dim room. Similarly, Azinović et al. [2] additionally attach polarization foils and capture two such sequences with perpendicular polarization orientation to separate skin surface and subsurface reflectance. Using a co-located light and camera setup, these methods assume the position of the dominant light source is known and no shadowing term needs to be modeled. Instead of using a smartphone flashlight, SunStage [57] takes a selfie video rotating under the sun as input and uses the varying angles between the sun and the face as guidance. Cast shadows from the sun are modeled by shadow mapping. It also jointly optimizes the sun’s position together with face geometry and appearance. All these methods assume the specular reflection (and the shadowing) of the face comes from a single dominant point or directional light source in the capture environment, and the ambient light contributes only to a low-frequency diffuse term. NextFace [16, 17, 18] exploits a differentiable ray tracer to estimate various scene attributes given unconstrained monocular images. It relies heavily on statistical shape and appearance priors [7, 24, 54] which limits its expressiveness. Our work instead focuses on accurate personalized reconstruction. Rainer et al. [51] use tiny shading networks to disentangle shading from explicit reflectance maps but assume a known high-quality geometry and smooth lighting. NeuFace [68] represents the face with a neural SDF and proposes to learn appearance factorization under unknown low frequency light with a novel neural BRDF basis. However, it requires multi-view input similar to those from a light stage setup. Closest to ours, FLARE [5] builds relightable head avatars from monocular videos. It uses pre-filtered environment maps for relighting and a neural approximation of it during training. This rendering model ignores the self-occlusion and bakes part of the shading into the albedos. In contrast, we explicitly handle light visibility, using ray tracing for diffuse reflectance and visibility modulated pre-filtered environment map for the specular reflectance, leading to higher quality shape and appearance reconstruction.
@@ 3 Monocular Appearance Capture
We now describe our method for facial appearance capture given a monocular head rotation sequence. We assume the expression of the subject does not change throughout the sequence. As a pre-processing step, we run monocular tracking based on landmarks [12] and a photometric loss [50] to obtain an initial canonical mesh from a 3DMM fit. We also estimate a fixed camera pose, per-frame rigid head poses and neck rotation. For more details about the pre-processing step please see the supplemental PDF. The output of our inverse rendering system will be a 3D asset of the subject containing a high-quality mesh, and diffuse albedo, specular intensity and roughness as 2D texture maps. In the following, we first describe our geometry optimization formulation in Section 3.1, shading model in Section 3.2, and then implementation details in Section 3.3.
@@ 3.1 Geometry Optimization
Balancing the updates to both geometry and textures in inverse rendering can be a challenging task. More specifically, the optimization might overfit too quickly on the textural components before learning the correct geometry. This is often the case when Laplacian shape regularization is applied to enforce geometric smoothness, making the geometry update too slow. Related work such as FLARE [5] addresses this issue by using a two-stage approach. First, a detailed geometry and only blurry textures are learned, and then, the shape is fixed and textures are learned after re-initialization in the second stage. Our method, however, optimizes geometry and textures at the same time. To do so, we adopt a preconditioning framework similar to Nicolet et al. [48], which biases gradient steps towards smooth solutions. The vertex positions $\mathbf{v}$ in each iteration are updated by
@@
3.2 Shading Model
Following the rendering equation [30], we compute the outgoing radiance $L(\omega_{o})$ in direction $\omega_{o}$ by:
[…]
We decompose the BRDF $f(\mathbf{x},\omega_{i},\omega_{o})$ into a diffuse term $f_{d}$ and a specular term $f_{s}$. We use the simple Lambertian model for the diffuse term:
@@ 3.2 Shading Model
where $\rho$ is the diffuse albedo. We use a specular BRDF similar to Kelemen and Szirmay-Kalos [34], which has been shown to be well suited for rendering human skin [47, Chapter 14]:
@@
Accounting for light visibility.
We consider direct illumination where $L_{i}(\omega_{i})$ comes only from the light sources. Spherical Harmonics [42, 13] or Spherical Gaussians [63] are popular representations used by prior methods [65, 68] but they can only model low- to medium-frequency lighting. Other work [5, 46] uses pre-filtered environment maps [32, 43], which allows to capture all-frequency lighting but still ignores self-shadowing. We aim to use a more realistic rendering model that accounts for light visibility to achieve higher photorealism. For the diffuse term, we apply explicit ray tracing to compute the outgoing radiance and use
[…]
where $V(\mathbf{x},\omega_{i})$ is the visibility at surface point $\mathbf{x}$ from incoming light direction $\omega_{i}$. We perform Monte Carlo integration with multiple importance sampling similar to Hasselgren et al. [28] skipping the denoising step as we did not observe any noise issue for the diffuse term in our experiments.
We empirically found that ray tracing for the specular term tends to produce flickery renders (Figure 6) due to our challenging scenario where only one camera view is available. Therefore, we propose to use a pre-filtered environment map but modulate the specular render with a view-dependent visibility term. We apply a split-sum approximation similar to Karis [31] and use
$L_{\text{specular}}(\omega_{o})\approx$ (6)
[…]
The first term is the material integral that depends on $\omega_{i}\cdot\mathbf{n}$ and the roughness. The second integral is the pre-filtered environment map. Both integrals can be pre-computed and stored as 2D textures. At rendering time, the rendering integral can be evaluated using only two texture lookups. When light visibility is not accounted for, the second integral only depends on the reflected view direction $\omega_{r}$ and the roughness $r$. When $r\ll 1$, the second integral is 0 unless $\omega_{i}$ is close to $\omega_{r}$. Therefore, we can approximate it as
@@
[…]
Here, $\tilde{V}(\mathbf{x},\omega_{r})$ is the view-dependent specular visibility. The simplest choice of $\tilde{V}(\mathbf{x},\omega_{r})$ is to only evaluate the light visibility at $\omega_{r}$, i.e., $\tilde{V}(\mathbf{x},\omega_{r})\coloneqq V(\mathbf{x},\omega_{r})$. Intuitively, this means the specular component is 0 if the reflected view direction is occluded. Note that this approximation is exact for perfect specular reflection (i.e. mirrors). To add some softness in the visibility attenuation term (instead of considering it as a binary function), we approximate it by Monte Carlo integration,
$\tilde{V}(\mathbf{x},\omega_{r})\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{V(\mathbf{x},\omega_{k})}{p(\omega_{k})},$ (8)
where the samples are drawn following the normal distribution of the specular lobe. This bears some resemblance to the appearance models of Saito et al. [53]. However, they use Spherical Gaussians for the specular lobe and parameterize the view-dependent specular visibility term using a neural network as they work with a large amount of studio data where the lighting is controlled and calibrated.
3.3 Implementation Details
Our training loss consists of an $L_{1}$ data term $\mathcal{L}_{\text{img}}$ and some regularization terms. We employ an $L_{1}$ mask loss $\mathcal{L}_{\text{mask}}$ between the mask obtained from MODNet [33] and the predicted binary mask. Although we use a parameterization similar to Nicolet et al. [48], we find it helpful to still employ a Laplacian regularizer to stabilize the geometry optimization such that we do not need to set different $\lambda_{\text{geo}}$ for different datasets. This regularizer encourages the Laplacian of the optimized mesh to stay close to the Laplacian of the initial 3DMM fit
@@ 3.3 Implementation Details
$\lambda_{\text{light}}\mathcal{L}_{\text{light}}+\lambda_{\text{rough}}\mathcal{L}_{\text{rough}}+\lambda_{\text{diffuse}}\mathcal{L}_{\text{diffuse}},$
where the weights are defined as
$\lambda_{\text{mask}}=0.1,\lambda_{\text{Lap}}=10,\lambda_{\text{light}}=0.1,$ (12)
$\lambda_{\text{rough}}=0.1,\lambda_{\text{diffuse}}=0.01.$
We capture 500 to 800 frames for each subject. Please see the supplemental PDF for details of our capture protocol. We then uniformly sample around 250 frames to use in the inverse rendering. We did not observe a performance gain or drop when using all of the frames. The images are cropped to 1K resolution. We also solve the texture maps at 1K resolution. The environment map is a cubemap with resolution $6\times 256\times 256$ at the largest mip-level, and resolution $6\times 8\times 8$ at the smallest mip-level. For each pixel, we draw $256$ light samples and $256$ BRDF (cosine) samples for the diffuse render, and $64$ samples to estimate the view-dependent specular visibility. We employ the Adam optimizer [19] with a learning of $0.1$ for vertex positions and the environment map, and $0.001$ for the textures. We use the differentiable rasterizer from Laine et al. [36] to obtain the primary visibility and the OptiX [49] engine for ray tracing. Each subject is trained for $6000$ iterations which takes around 2 hours on a Nvidia RTX 3090 GPU.
@@ 4.1 Appearance and Geometry Reconstruction
We begin by highlighting the versatility of our approach by showing several appearance capture results of different subjects in different environments in Fig. 1. Each row of the figure illustrates one of the input images, the corresponding render using our recovered appearance properties, and then a breakdown of the reconstructed appearance maps (diffuse albedo, specular intensity, specular roughness) and geometry. Our method can be applied indoors or outdoors, with sunny or cloudy skies. We show how the recovered geometry and appearance maps can be used to render the subject in a new environment by relighting them. Fig. 2 shows two subjects relit in two different environments (see Fig. 1 for the original environment).

### v2

@@ 1 Introduction
Like facial geometry reconstruction, the field of skin appearance estimation is also well-studied in controlled studio environments, where accurate appearance maps can be reconstructed from large setups that obtain multi-view or multi-shot images under calibrated lighting [55]. Following the geometry trend, current research aims to allow facial appearance capture in less constrained settings, for example outdoors using the sun as a single point light [62]. Unfortunately, these methods often make simplifying assumptions, and thus there still exists a large gap in reconstruction quality between current in-the-wild methods and production-ready studio-based capture.
In this work we present a new method for facial appearance capture in the wild, surpassing the level of fidelity of existing lightweight methods. Our approach requires only a short video sequence of a simple head rotation, captured from a single camera in any environment, including indoors or outdoors, on a sunny day or in shadow. Our approach is built on traditional inverse rendering optimization, where a fast differentiable renderer is used to solve for the geometry and appearance parameters together with the environment lighting simultaneously. Different from previous methods, we do not make any assumptions on the lighting condition (e.g. we do not require a sun in the sky), and as our main contribution we explicitly take visibility into account, effectively removing baked-in shading by correctly modeling self-occlusion in our appearance solver. The result is a detailed geometry mesh with textures for diffuse albedo, specular intensity and roughness. As we will show, our approach leads to more faithful recovery of the appearance properties than existing techniques in the wild. As a particular application, our approach allows fast capture of actors on a film set, with resulting assets that can be used directly in traditional VFX pipelines. In summary, we make the following key contributions:
@@
  • A novel shading model that explicitly handles visibility and self-occlusion for inverse rendering pipelines, achieving high-quality appearance reconstruction from monocular input.
@@ 2 Related Work
Inverse rendering is the process of decomposing the scene into 3D shape, material and illumination by simulating the rendering process and comparing the results against captured images. It has been a popular research topic with the recent advances in novel view synthesis using neural implicit representations [46, 47, 61], 3D Gaussian splatting [36], mesh-based differentiable renderers (rasterizers [37] and path tracers [40, 30]) and using diffusion models as prior [43, 21, 41, 65]. While some existing techniques target a very challenging scenario where the lighting can differ across different images [8, 9, 20], we restrict our discussion here to methods that assume a static unknown environment lighting. PhySG [70] utilizes spherical Gaussians to approximately and efficiently evaluate the rendering equation in closed form. Munkberg et al[48] propose an efficient end-to-end framework for joint learning of topology, triangle meshes and materials, achieving much faster training and inference compared to previous NeRF-based factorization methods [72, 7, 59]. They also introduce a differentiable formulation of the split-sum approximation of environment lighting to efficiently recover all-frequency lighting. Follow-up work [29] shows that material and lighting decomposition can be further improved with a more realistic shading model, incorporating ray tracing and Monte Carlo integration. Recently, 3D Gaussian Splatting techniques were used in conjunction with physical-based rendering to allow for scene relighting [42, 23, 5, 74, 64].
@@
Traditional face capture studios often employ a multi-view setup with controlled and calibrated lighting conditions to reconstruct the skin appearance [55, 25, 26, 14]. Similar setups were used for facial appearance decomposition with neural and gaussian primitives [56, 69, 67, 66, 57]. Lighter alternative setups were also explored by Lattas et al[39] and Choi et al[13] but they still require the subject to be seating in a dedicated space. Recently, the research community has investigated more lightweight setups which are easily accessible to everyone. However, most of these techniques still pose some constraints on the capture environment. CoRA [28] reconstructs relightable 3D face assets from a single co-located smartphone flashlight sequence captured in a dim room. Similarly, Azinović et al[1] additionally attach polarization foils and capture two such sequences with perpendicular polarization orientation to separate skin surface and subsurface reflectance. Using a co-located light and camera setup, these methods assume the position of the dominant light source is known and no shadowing term needs to be modeled. Instead of using a smartphone flashlight, SunStage [62] takes a selfie video rotating under the sun as input and uses the varying angles between the sun and the face as guidance. Cast shadows from the sun are modeled by shadow mapping. It also jointly optimizes the sun’s position together with face geometry and appearance. All these methods assume the specular reflection (and the shadowing) of the face comes from a single dominant point or directional light source in the capture environment, and the ambient light contributes only to a low-frequency diffuse term. Another line of work [15, 16, 17, 38, 27, 51, 18] aims to reconstruct shape and reflectance properties from a single portrait image, often relying heavily on statistical shape and appearance priors [6, 24, 58] which limits its expressiveness. Our work instead focuses on accurate personalized reconstruction. Rainer et al[54] use tiny shading networks to disentangle shading from explicit reflectance maps but assume a known high-quality geometry and smooth lighting. NeuFace [73] represents the face with a neural SDF and proposes to learn appearance factorization under unknown low frequency light with a novel neural BRDF basis. However, it requires multi-view input similar to those from a light stage setup. Closest to ours, FLARE [4] builds relightable head avatars from monocular videos. It adopts the split-sum approximation [32, 48] for relighting and a neural version of it during training. This rendering model ignores the self-occlusion and bakes part of the shading into the albedos. In contrast, we propose a modified formulation of the split-sum approximation which explicitly handles light visibility and combine it with ray tracing, leading to higher quality shape and appearance reconstruction.
@@ 3 Monocular Appearance Capture
We now describe our method for facial appearance capture given a monocular head rotation sequence. We assume the expression of the subject does not change throughout the sequence. As a pre-processing step, we run monocular tracking based on landmarks [11] and a photometric loss [53] to obtain an initial canonical mesh from a 3DMM fit. We also estimate a fixed camera pose, per-frame rigid head poses and neck rotation. For more details about the pre-processing step please see the supplemental PDF. The output of our inverse rendering system will be a 3D asset of the subject containing a high-quality mesh, and diffuse albedo, specular intensity and roughness as 2D texture maps. In the following, we first describe our geometry optimization formulation in Section 3.1, our novel occlusion-aware shading model in Section 3.2, and then optimization details in Section 3.3.
@@ 3.1 Geometry Optimization
Balancing the updates to both geometry and textures in inverse rendering can be a challenging task. More specifically, the optimization might overfit too quickly on the textural components before learning the correct geometry. This is often the case when Laplacian shape regularization is applied to enforce geometric smoothness, making the geometry update too slow. Related work such as FLARE [4] addresses this issue by using a two-stage approach. First, a detailed geometry and only blurry textures are learned, and then, the shape is fixed and textures are learned after re-initialization in the second stage. Our method, however, optimizes geometry and textures at the same time. To do so, we adopt a preconditioning framework similar to Nicolet et al[50], which biases gradient steps towards smooth solutions. The vertex positions $\mathbf{v}$ in each iteration are updated by
@@
3.2 Occlusion-Aware Shading Model
Following the rendering equation [31], we compute the outgoing radiance $L(\omega_{o})$ at location $\mathbf{x}$ from direction $\omega_{o}$ by:
$L(\omega_{o})=\int_{\Omega}f(\mathbf{x},\omega_{i},\omega_{o})L_{i}(\omega_{i})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$ (2)
We decompose the BRDF $f(\mathbf{x},\omega_{i},\omega_{o})$ into the sum of a diffuse term $f_{d}$ and a specular term $f_{s}$. We use the simple Lambertian model for the diffuse term:
@@ 3.2 Occlusion-Aware Shading Model
where $\rho$ is the diffuse albedo. We use a specular BRDF similar to Kelemen and Szirmay-Kalos [35], which has been shown to be well suited for rendering human skin [49, Chapter 14]:
@@
Accounting for self-occlusion.
We consider direct illumination where $L_{i}(\omega_{i})$ comes only from the light sources. Spherical Harmonics [44, 12] or Spherical Gaussians [68] are popular representations used by prior methods [70, 73] but they can only model low- to medium-frequency lighting. We therefore follow other work, e.g. FLARE [4], and use a differentiable split-sum approximation [33, 45, 48], which allows to capture all-frequency lighting. Unfortunately this approach ignores self-shadowing, and so we propose a novel visibility-modulated split-sum approximation to account for self-shadowing, which we describe in the following.
We start by introducing the split-sum formulation from Karis [32] which approximates  Eq. 2 as
$L(\omega_{o})\approx$ (5)
$\int_{\Omega}f(\mathbf{x},\omega_{i},\omega_{o})(\omega_{i}\cdot\mathbf{n})d\omega_{i}\int_{\Omega}L_{i}(\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$
The first term is the integral of the BRDF under a solid white environment map, i.e. $L_{i}(\omega_{i})=1,\forall\omega_{i}$. It only depends on $\cos\theta=\omega_{i}\cdot\mathbf{n}$ and the roughness $r$. Therefore, it can be precomputed and stored as a 2D look-up texture. The second integral is a pre-filtered environment map, where $D$ is the normal distribution function of the BRDF and $\mathbf{h}=\frac{\omega_{i}+\omega_{o}}{\|\omega_{i}+\omega_{o}\|_{2}}$ is the half vector. This term can also be precomputed as a mipmap by convolving the environment map with $D$ at different roughness values. At rendering time, the outgoing radiance can then be efficient computed using only two texture lookups. Note that when evaluating  Eq. 5, the contribution of the prefiltered environment map is only dependent on the reflected view direction $\omega_{r}$ and the roughness $r$. This is not physically accurate as the contribution should be reduced if the light is (partially) blocked from the upper hemisphere at a specific point. To account for self-occlusion, we must modulate $L_{i}(\omega_{i})$ differently at different locations based on the light visibility $V(\mathbf{x},\omega_{i})$, i.e., making the second integral in  Eq. 5 $\int_{\Omega}L_{i}(\omega_{i})V(\mathbf{x},\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}$. This new integral cannot be precomputed anymore since it’s different for different $\mathbf{x}$. However, we notice that when $r\ll 1$, the new integral is 0 unless $\omega_{i}$ is close to $\omega_{r}$. Therefore, we can approximate it as
@@
$\int_{\Omega}L_{i}(\omega_{i})V(\mathbf{x},\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}$ (6)
$\approx$ $\tilde{V}(\mathbf{x},\omega_{r})\int_{\Omega}L_{i}(\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$
Here, $\tilde{V}(\mathbf{x},\omega_{r})$ is the view-dependent visibility. The simplest choice of $\tilde{V}(\mathbf{x},\omega_{r})$ is to only evaluate the light visibility at $\omega_{r}$, i.e., $\tilde{V}(\mathbf{x},\omega_{r})\coloneqq V(\mathbf{x},\omega_{r})$. Intuitively, this means the outgoing radiance is 0 if the reflected view direction is occluded. Note that this approximation is exact for perfect specular reflection (i.e. mirrors). To add some softness in the visibility term (instead of considering it as a binary function), we approximate it by Monte Carlo integration,
$\tilde{V}(\mathbf{x},\omega_{r})\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{V(\mathbf{x},\omega_{k})}{D(\mathbf{n},\omega_{k},\omega_{r},r)},$ (7)
where the samples are drawn following the normal distribution of the BRDF. This bears some resemblance to the appearance models of Saito et al[56]. However, they parameterize the view-dependent specular visibility term using a neural network as they work with a large amount of studio data where the lighting is controlled and calibrated. Note however,  Eq. 6 introduces large errors for rough surfaces (i.e., when $r\ll 1$ does not hold). We therefore use it only for the specular component, and implemented ray tracing with multiple importance sampling [60] for the diffuse component using the OptiX [52] engine.
3.3 Optimization Details
Our reconstruction loss consists of an $L_{1}$ data term $\mathcal{L}_{\text{img}}$ and some regularization terms. We employ an $L_{1}$ mask loss $\mathcal{L}_{\text{mask}}$ between the mask obtained from MODNet [34] and the predicted binary mask. Although we use a parameterization similar to Nicolet et al[50], we find it helpful to still employ a Laplacian regularizer to stabilize the geometry optimization such that we do not need to set different $\lambda_{\text{geo}}$ for different datasets. This regularizer encourages the Laplacian of the optimized mesh to stay close to the Laplacian of the initial 3DMM fit
@@ 3.3 Optimization Details
$\lambda_{\text{light}}\mathcal{L}_{\text{light}}+\lambda_{\text{rough}}\mathcal{L}_{\text{rough}}+\lambda_{\text{diffuse}}\mathcal{L}_{\text{diffuse}}.$
@@ 4.1 Appearance and Geometry Reconstruction
We begin by highlighting the versatility of our approach by showing several appearance capture results of different subjects in different environments in Fig. 1. Each row of the figure illustrates one of the input images, the corresponding render using our recovered appearance properties, and then a breakdown of the reconstructed appearance maps (diffuse albedo, specular intensity, specular roughness) and geometry. Our method can be applied indoors or outdoors, with sunny or cloudy skies. We show how the recovered geometry and appearance maps can be used to render the subject in a new environment by relighting them. Fig. 2 shows two subjects relit in two different environments (see Fig. 1 for the original environment). Please refer to the supplemental video for more results.